# Continuous-variable quantum key distribution system: A review and perspective

### v2

@@ IV.1.3 Field tests
IV.1.4 System on chip
Photonic integrated circuit is a promising way to realize a large scale and cost effective system. After the design is finalized, the production cost of the chip will sharply decrease with the increase of production. Therefore, in addition to miniaturizing the system, chipization can also greatly promote the low-cost mass production for the cost sensitive CV-QKD system. The earliest attempt to chip-based CV-QKD systems is in 2015, where a silicon photonic chip comprising all major CV-QKD components as well as complete subsystems is designed and fabricated Ziebell et al. (2015).
@@ IV.1.4 System on chip
Figure 29: The first chip-based CV-QKD system. From G. Zhang et al. Zhang et al. (2019b). Reproduced with permission from Nat. Photonics 13, 839 (2019). Copyright 2019 Springer Nature Limited.
@@
This work demonstrate that the Silicon-On-Insulator platform can basically satisfy the requirement of a CV-QKD system. A two-dimensional grating coupler integrated on chip can be directly used for the polarization multiplexing and de-multiplexing. The pulse generation, the Gaussian modulation, the variable optical attenuator and the detector are all realized with chip-based components. The insertion loss of the grating coupler, the detection efficiency and the bandwidth of the chip-based detector are the concerns, where a part of these issues are solved in the later works, and we detailed them in the local LO system part. A comparison of the secret key rate of different systems is shown in Fig. 30.
@@ IV.1.5 Other in-line LO systems
Figure 31: The in-line LO CV-QKD systems with particular modulation format. (a) The 4-state modulated in-line LO CV-QKD system Hirano et al. (2017), from T. Hirano et al. (b) The unidimensional modulated in-line LO CV-QKD system Wang et al. (2017b), from X. Wang et al. (a) Reproduced with permission from Quantum Sci. Technol. 2, 024010 (2017). Copyright 2017 IOP Publishing Ltd. (b) Reproduced with permission from Phys. Rev. A 95, 062330 (2017). Copyright 2017 American Physical Society.
@@
2020 Gaussian AM+PM Frequency and polarization 100 MHz 25 km 1.85 Mbps Wang et al. (2020a)
2022 QPSK IQ modulator Frequency and polarization 5 GBaud 25 km 2.48 Mbps Wang et al. (2022b)
@@ IV.2.4 Recent progress
A recent work with the structure shown in Fig. 34 (c) can realize a transmission distance over 100 km, and the asymptotic secret key rate can reach 10.36 Mbps, 2.59 Mbps and 0.69 Mbps over transmission distance of 50 km, 75 km and 100 km Pi et al. (2023). This work significantly increases the transmission distances of a local LO CV-QKD system, and shows a promising way to realize long-distance and high-speed QKD using telecom devices.
@@ IV.3 CV-QKD systems co-existed with classical communication environment
The homodyne and heterodyne detection in CV-QKD system act as a matched filter since the LO naturally introduce a frequency selection on the received signal, therefore the filter in time and frequency domain is unnecessary. Moreover, the devices in a CV-QKD system is compatible with the classical coherent optical communications. Therefore, CV-QKD is suitable for coexisting with classical signals, which is easy for deployments.
@@ IV.4.1 Free space systems
In 2019, the effective resistance against background noise of CV-QKD is theoretically and experimentally demonstrated Wang et al. (2019f). In 2020, as shown in Fig. 36 (b), a phase compensation strategy is developed and experimental demonstrated Wang et al. (2020b). This work checks the correlation between data held by Alice and Bob in a free space channel, and proves that the fluctuation of transmittance disappears in the correlation, thus enabling phase compensation for signals over fluctuant channels. Later, a series of methods for free space polarization compensation Wang et al. (2020c), data and transmittance synchronization Wang et al. (2021a), data acquisition Wei et al. (2023) are proposed, and the feasibility of secure key distribution with free space CV-QKD through fog is experimentally demonstrated Wang et al. (2021b). Recently, a passive-state-preparation CV-QKD system shown in Fig. 36 (c) with free space channel is experimentally demonstrated Zhang et al. (2023). Thermal-state polarization multiplexing transmitted LO, synchronized channel transmittance monitoring and fine-grained phase compensation techniques are proposed to support a secret key rate of 1.015 Mbps with a free space channel of -15 dB simulated transmittance.
@@ V THE ADVANCED CV-QKD SYSTEM PROGRESS FOR FUTURE APPLICATIONS
V.1 Digital CV-QKD system
Impairment compensation on digital domain can significantly simplify the system structure, contributing to a simple and stable system. The study on classical coherent communications promotes the DSP algorithms in a CV-QKD system. For instance, at the transmitter side, the pulse shaping algorithm is used to raise the availability of the frequency band. For the receiver, the frequency shift is recognized and the down conversion is completed in digital domain, time recovery algorithm is used to obtain the optimal sampling points, digital filter is used to reduce the spectrum mismatch between the transmitter and receiver site, and various of algorithms are developed to distill the parameters of polarization compensation and phase compensation. Thanks to the wide application of DSP in CV-QKD systems, the speed of the system is becoming higher and higher, the latest achievement reach the baud rate of 10 GBaud Hajomer et al. (2023a).
@@ V.1 Digital CV-QKD system
Figure 38: The local LO CV-QKD systems with chip-based devices. (a) A silicon based receiver of the CV-QKD system. From Y. Pietri et al. Piétri et al. (2023). (b) A InP transmitter of the CV-QKD system. From J. Aldama et al. Aldama et al. (2023b). (c) The 10 GBaud high-speed CV-QKD system with chip-based receiver. From A. Hajomer et al. Hajomer et al. (2023a). (a) Reproduced with permission from OFC. M1I.2 (2023). Copyright 2023 The authors. (b) Reproduced with permission from OFC. M1I.3 (2023). Copyright 2023 The authors. (c) Reproduced with permission from arXiv 2305.19642 (2023). Copyright 2023 arXiv.
@@
V.2 Chip-based local LO system
The advanced local LO CV-QKD system is developing towards compact module with photonics integration, benifiting from stability and scalability, which enables cost-effective deployments in large-scale. Recently, a seires of investigations have been carried out on high-performance chip-based transmitter and receiver for local LO CV-QKD system Wang et al. (2020d); Li et al. (2021b); Wang et al. (2022d); Aldama et al. (2023b); Piétri et al. (2023); Hajomer et al. (2023a); Li et al. (2023b, b); Jia et al. (2023); Luo et al. (2023). There are two mainstream fabrication platforms, Silicon-On-Insulator and III-V.
Figure 40: The secret key rate of a single user (solid line) and the overall network (dashed line) with different number of accessed users.
The Silicon-On-Insulator platform has the advantages of low cost and good ductility, which can utilize mature silicon CMOS processes to manufacture optical devices. The refractive index of the silicon waveguide is 3.42, which can form a significant refractive index difference with silicon dioxide, ensuring that the silicon waveguide can have a smaller waveguide bending radius, which is beneficial for high-density device integration. Recently, a Silicon-On-Insulator CV-QKD receiver shown in Fig. 38 (a) for a digital system is tested. The bandwidth is significantly enhanced to 250 MHz which makes detect the frequency division multiplexed quantum and pilot signal possible Piétri et al. (2023). The maximal detection efficiency of the overall receiver is 0.26, with a shot noise-to-electronic-noise ratio of 20 dB at low frequencies and more than 7 dB for 250 MHz. Under the untrusted loss of 1.38 dB, a secret key rate of 280 kbps is achieved with excess noise of 0.1102 at Alice’s side.
@@ V.2 Chip-based local LO system
One promising solution is using III-V platform such as InP, the other way is making heterogeneous integration, where the III-V chip-based laser is combined with the silicon chip via bonding or growing. Some recent researches have shown the feasibility of implementing the above two technical routines. J. Aldama et al. has demonstrated a InP CV-QKD transmitter Aldama et al. (2023b) shown in Fig. 38 (b), including an electro-absorption modulator, an IQ modulator and a variable optical attenuator. These three devices are cascaded and integrated on one chip, supporting more than 1 GHz bandwidth. 0.4 and 2.3 Mbps secret key rate is achieved via a test with 11 km fiber and back-to-back connection, which verfies the possibility of using InP platform to produce a transmitter satisfying the requirement of CV-QKD. Once the laser diode is integrated to the InP chip, a fully integrated CV-QKD chip can be realized.
Recently, using the silicon based on-chip receiver with high-efficiency Ge photodiodes, a high-speed system at 10 GBaud is realized Hajomer et al. (2023a), as shown in Fig. 38 (c). The system is able to generate high secret key rates exceeding 0.7 Gbps over a distance of 5km and 0.3 Gbps over a distance of 10km, paving the way of the high-performance and compact CV-QKD system.
@@
V.3 Point-to-multipoint system
@@
A high-rate CV-QKD access network is realized, as shown in Fig. 41. For the transmitter, quantum signal is generated with an IQ modulator, which is multiplexed with pilot tone signal with frequency division multiplexing and polarization multiplexing. For the receiver, the de-multiplexing is realized by a polarization controller and a polarization beamsplitter. Then, quantum signal and pilot tone are seperately detected by heterodyne detection. The average secret key rate of each user can reach 4.1 Mbps at 15 km when the network capacity is 4, with the repetition frequency of 500 MHz Pan et al. . Further, the capacity of the network is extended to 8, and the secret key rate is 7.44 Mbps at 6 km for each user Bian et al. (2023b). This result can well support the quantum access network, such as linking the end users within a campus (see Fig. 42).
@@ VI.1 Attacks and countermeasures
The local LO system is proposed to defend the attack at the LO by generating LO locally. Moreover, since the classical pilot tone is transmitted in the unsecured channel, some attacking schemes at the pilot tone are proposed such as the reference pulse attack and so on Ren et al. (2019); Shao et al. (2022); Huang, Huang, and Peng (2019).
@@
VI.2 Measurement-device-independent system
@@
Figure 43: The experimental schemes of the CV-MDI QKD. (a) The first experimental demonstration. From S. Pirandola et al. Pirandola et al. (2015). (b) The first CV-MDI QKD with optical fiber links. From Y. Tian et al. Tian et al. (2022). (c) The first experimental demonstration without locking systems. From A. Hajomer et al. Hajomer et al. (2023c). (a) Reproduced with permission from Nat. Photonics 9, 397 (2015). Copyright 2015 Nature Publishing Group. (b) Reproduced with permission from Optica 9, 492 (2022). Copyright 2022 Optica Publishing Group. (c) Reproduced with permission from OFC. M2I.2 (2023). Copyright 2023 The Authors.
@@
In 2023, the frequency and optical phase locking in a CV-MDI QKD system is removed by using a new relay structure based on a polarization-based 90-degree optical hybrid and a well-designed DSP pipeline Hajomer et al. (2023c), shown in Fig. 43 (c). Moreover, this system uses continuous-wave laser with digital pulse shaping and digital time synchronization, therefore the additional amplitude modulation for pulse generation and a delay line for time synchronization is unneccessary. A secret key rate of 600 kbps over 2 dB loss channel is achieved.
@@ VII SUMMARY AND OUTLOOK
There is still the need to develop and design high-performance protocol, including the deep exploration on the insight between shot-noise-unit and security, and developing novel protocols with the secret key rate exceeding Gaussian protocols under same launch energy, which further approaching the fundamental rate-distance limit of point-to-point quantum key distribution while still keeps simple system structure. In practice, different secret key rates might be considered, for instance, with respect to individual, collective, or fully coherent attacks. The choice of these rates may also be associated with a specific sublevel of security to be reached.
@@ Acknowledgements.
We are thankful for the enlightening discussions with and helpful comments from reviewers and numerous colleagues, including R. Goncharov, P. Huang, L. Huang, Y. Li, Q. Liao, T. Matsuura, S. Pirandola, S. Sarmiento, A. Vidiella-Barranco, T. Wang, X. Wang and B. Xu. This work was supported by the National Natural Science Foundation of China (62001044, 62201013), the Basic Research Program of China (JCKY2021210B059), the Equipment Advance Research Field Foundation (315067206), and the Fund of State Key Laboratory of Information Photonics and Optical Communications (IPOC2021ZT02).

### v3

@@ IV.1.3 Field tests
IV.1.4 Systems on chip
Figure 29: The first chip-based CV-QKD system. From G. Zhang et al. Zhang et al. (2019b). Reproduced with permission from Nat. Photonics 13, 839 (2019). Copyright 2019 Springer Nature.
Photonic integrated circuit is a promising way to realize a large scale and cost effective system. After the design is finalized, the production cost of the chip will sharply decrease with the increase of production. Therefore, in addition to miniaturizing the system, chipization can also greatly promote the low-cost mass production for the cost sensitive CV-QKD system. The earliest attempt to chip-based CV-QKD systems is in 2015, where a silicon photonic chip comprising all major CV-QKD components as well as complete subsystems are designed and fabricated Ziebell et al. (2015).
@@ IV.1.4 Systems on chip
This work demonstrates that the Silicon-On-Insulator platform can basically satisfy the requirement of a CV-QKD system. A two-dimensional grating coupler integrated on chip can be directly used for the polarization multiplexing and de-multiplexing. The pulse generation, the Gaussian modulation, the variable optical attenuator and the detector are all realized with chip-based components. The insertion loss of the grating coupler, the detection efficiency and the bandwidth of the chip-based detector are the concerns, where a part of these issues are solved in the later works, and we detailed them in the local LO system part. A comparison of the secret key rate of different systems is shown in Fig. 30.
@@ IV.1.5 Other in-line LO systems
Figure 31: The in-line LO CV-QKD systems with particular modulation format. (a) The 4-state modulated in-line LO CV-QKD system Hirano et al. (2017), from T. Hirano et al. (b) The unidimensional modulated in-line LO CV-QKD system Wang et al. (2017b), from X. Wang et al. (a) T. Hirano, Quantum Sci. Technol., 2, 024010, 2017; licensed under a Creative Commons Attribution ($C$ BY) license. (b) Reproduced with permission from Phys. Rev. A 95, 062330 (2017). Copyright 2017 American Physical Society.
@@
2020 Gaussian AM+PM Frequency and polarization 100 MHz 25 km 7.04 Mbps Wang et al. (2020a)
2022 QPSK IQ modulator Frequency and polarization 5 GBaud 10 km 133.6 Mbps Wang et al. (2022b)
@@ IV.2.4 Recent progresses
A recent work with the structure shown in Fig. 34 (c) can realize a transmission distance over 100 km, and the asymptotic secret key rate can reach 7.55 Mbps, 1.87 Mbps and 0.51 Mbps over transmission distance of 50 km, 75 km and 100 km Pi et al. (2023). This work significantly increases the transmission distances of a local LO CV-QKD system, and shows a promising way to realize long-distance and high-speed QKD using telecom devices.
@@ IV.3 CV-QKD systems co-existed with classical communication environment
The homodyne and heterodyne detection in CV-QKD system act as a matched filter since the LO naturally introduces a frequency selection on the received signal, therefore the filter in time and frequency domain is unnecessary. Moreover, the devices in a CV-QKD system is compatible with the classical coherent optical communications. Therefore, CV-QKD is suitable for coexisting with classical signals, which is easy for deployments.
@@ IV.4.1 Free space systems
In 2019, the effective resistance against background noise of CV-QKD is theoretically and experimentally demonstrated Wang et al. (2019e). In 2020, as shown in Fig. 36 (b), a phase compensation strategy is developed and experimental demonstrated Wang et al. (2020b). This work checks the correlation between data held by Alice and Bob in a free space channel, and proves that the fluctuation of transmittance disappears in the correlation, thus enabling phase compensation for signals over fluctuant channels. Later, a series of methods for free space polarization compensation Wang et al. (2020c), data and transmittance synchronization Wang et al. (2021a), data acquisition Wei et al. (2023) are proposed, and the feasibility of secure key distribution with free space CV-QKD through fog is experimentally demonstrated Wang et al. (2021b). Recently, a passive-state-preparation CV-QKD system shown in Fig. 36 (c) with free space channel is experimentally demonstrated Zhang et al. (2023). Thermal-state polarization multiplexing transmitted LO, synchronized channel transmittance monitoring and fine-grained phase compensation techniques are proposed to support a secret key rate of 1.015 Mbps with a free space channel of -15 dB simulated transmittance.
@@ V THE ADVANCED CV-QKD SYSTEM PROGRESS FOR FUTURE APPLICATIONS
V.1 Digital CV-QKD systems
Impairment compensation on digital domain can significantly simplify the system structure, contributing to a simple and stable system. The study on classical coherent communications promotes the DSP algorithms in a CV-QKD system. For instance, at the transmitter side, the pulse shaping algorithm is used to raise the availability of the frequency band. For the receiver, the frequency shift is recognized and the down conversion is completed in digital domain, time recovery algorithm is used to obtain the optimal sampling points, digital filter is used to reduce the spectrum mismatch between the transmitter and receiver site, and various of algorithms are developed to distill the parameters of polarization compensation and phase compensation. Thanks to the wide application of DSP in CV-QKD systems, the speed of the system is becoming higher and higher, the latest achievement reaches the baud rate of 10 GBaud Hajomer et al. (2023a).
@@ V.1 Digital CV-QKD systems
Figure 38: The local LO CV-QKD systems with chip-based devices. (a) A silicon based high-speed CV-QKD receiver. From Y. Bian et al. Bian et al. (2024). (b) The 10 GBaud high-speed CV-QKD system with chip-based receiver. From A. Hajomer et al. Hajomer et al. (2023a). (b) Reproduced with permission from arXiv 2305.19642 (2023). Copyright 2023 The Authors.
@@
V.2 Chip-based local LO systems
The advanced local LO CV-QKD system is developing towards compact module with photonics integration, benifiting from stability and scalability, which enables cost-effective deployments in large-scale. Recently, a seires of investigations have been carried out on high-performance chip-based transmitter and receiver for local LO CV-QKD system Wang et al. (2020d); Li et al. (2021b); Wang et al. (2022d); Aldama et al. (2023b); Piétri et al. (2023); Hajomer et al. (2023a); Li et al. (2023b); Jia et al. (2023); Luo et al. (2023). There are two mainstream fabrication platforms, Silicon-On-Insulator and III-V.
The Silicon-On-Insulator platform has the advantages of low cost and good ductility, which can utilize mature silicon CMOS processes to manufacture optical devices. The refractive index of the silicon waveguide is 3.42, which can form a significant refractive index difference with silicon dioxide, ensuring that the silicon waveguide can have a smaller waveguide bending radius, which is beneficial for high-density device integration. In 2023, a Silicon-On-Insulator CV-QKD receiver for a digital system is tested. The bandwidth is significantly enhanced to 250 MHz which makes detect the frequency division multiplexed quantum and pilot signal possible Piétri et al. (2023). The maximal detection efficiency of the overall receiver is 0.26, with a shot noise-to-electronic-noise ratio of 20 dB at low frequencies and more than 7 dB for 250 MHz. Under the untrusted loss of 1.38 dB, a secret key rate of 280 kbps is achieved with excess noise of 0.1102 at Alice’s side. The photonics chip shown in Fig. 38 (a) is the CV-QKD receiver designed by Y. Bian et al., which achieves the bandwidth of over 2 GHz and the quantum to classical noise ratio of 5 dB at 2 GHz Bian et al. (2024). Using the silicon based on-chip receiver with high-efficiency Ge photodiodes, a high-speed system at 10 GBaud is realized Hajomer et al. (2023a), as shown in Fig. 38 (b). The system is able to generate high secret key rates exceeding 0.7 Gbps over a distance of 5 km and 0.3 Gbps over a distance of 10km, paving the way of the high-performance and compact CV-QKD system.
@@ V.2 Chip-based local LO systems
One promising solution is using III-V platform such as InP, the other way is making heterogeneous integration, where the III-V chip-based laser is combined with the silicon chip via bonding or growing. Some recent researches have shown the feasibility of implementing the above two technical routines. J. Aldama et al. has demonstrated a InP CV-QKD transmitter Aldama et al. (2023b) including an electro-absorption modulator, an IQ modulator and a variable optical attenuator. These three devices are cascaded and integrated on one chip, supporting more than 1 GHz bandwidth. 0.4 and 2.3 Mbps secret key rate is achieved via a test with 11 km fiber and back-to-back connection, which verfies the possibility of using InP platform to produce a transmitter satisfying the requirement of CV-QKD. Once the laser diode is integrated to the InP chip, a fully integrated CV-QKD chip can be realized.
@@
V.3 Point-to-multipoint systems
@@
Figure 40: Secret key rate versus transmission distance when the number of the Bobs is 8 (blue), 32 (orange) and 128 (green), including the secret key rate of a single Bob (solid line) and the secret key rate of the overall protocol (dashed line). The simulation parameters are as below, number of Bobs $N=8$, modulation variance $V_{M}=4$, reconciliation efficiency $\beta=95.6\%$, $\varepsilon_{tot}=0.0383$, the detection efficiency $\eta_{d}=60\%$, the detection noise $v_{el}=0.15$.
@@
A high-rate CV-QKD access network is realized, as shown in Fig. 41. For the transmitter, quantum signal is generated with an IQ modulator, which is multiplexed with pilot tone signal with frequency division multiplexing and polarization multiplexing. For the receiver, the de-multiplexing is realized by a polarization controller and a polarization beamsplitter. Then, quantum signal and pilot tone are seperately detected by heterodyne detection. The average secret key rate of each user can reach 4.1 Mbps at 15 km when the network capacity is 4, with the repetition frequency of 500 MHz Pan et al. (2023b). When the capacity of the network is extended to 8, the secret key rate for each user can reach 7.44 Mbps at 6 km Bian et al. (2023b), and 3.20 Mbps at 15 km Pan et al. (2024). This result can well support the quantum access network, such as linking the end users within a campus (see Fig. 42).
@@ VI.1 Attacks and countermeasures
The local LO system is proposed to defend the attack at the LO by generating LO locally. However, since the classical pilot tone is transmitted in the unsecured channel Shao et al. (2021), some attacking schemes at the pilot tone are proposed such as the reference pulse attack and so on Ren et al. (2019); Shao et al. (2022); Huang, Huang, and Peng (2019).
@@
VI.2 Measurement-device-independent systems
@@
Figure 43: The experimental schemes of the CV-MDI QKD. (a) The first experimental demonstration. From S. Pirandola et al. Pirandola et al. (2015). (b) The first CV-MDI QKD with optical fiber links. From Y. Tian et al. Tian et al. (2022). (c) The first experimental demonstration without locking systems. From A. Hajomer et al. Hajomer et al. (2023c). (a) Reproduced with permission from Nat. Photonics 9, 397 (2015). Copyright 2015 Springer Nature. (b) Reproduced with permission from Optica 9, 492 (2022). Copyright 2022 Optica Publishing Group. (c) Reproduced with permission from OFC. M2I.2 (2023). Copyright 2023 The Authors.
@@
In 2023, a simple and practical CV-MDI QKD system that can coexist with classical telecommunications channels is realized with a real-time frequency and phase locking system Hajomer, Andersen, and Gehring (2023). Later, the frequency and optical phase locking in a CV-MDI QKD system is removed by a new relay structure based on a polarization-based 90-degree optical hybrid and a well-designed DSP pipeline Hajomer et al. (2023c), shown in Fig. 43 (c). Moreover, this system uses continuous-wave laser with digital pulse shaping and digital time synchronization, therefore the additional amplitude modulation for pulse generation and a delay line for time synchronization is unneccessary. A secret key rate of 600 kbps over 2 dB loss channel is achieved.
@@ VII SUMMARY AND OUTLOOK
There is still the need to develop and design high-performance protocol, including the deep exploration on the insight between shot-noise-unit and security, and developing novel protocols with the secret key rate exceeding Gaussian protocols under same launch energy, which further approaches the fundamental rate-distance limit of point-to-point quantum key distribution while still keeps simple system structure. In practice, different secret key rates might be considered, for instance, with respect to individual, collective, or fully coherent attacks. The choice of these rates may also be associated with a specific sublevel of security to be reached.
@@ Acknowledgements.
We are thankful for the enlightening discussions with and helpful comments from reviewers and numerous colleagues, including R. Goncharov, A. Hajomer, P. Huang, L. Huang, Y. Li, Q. Liao, T. Matsuura, S. Pirandola, S. Sarmiento, A. Vidiella-Barranco, T. Wang, X. Wang and B. Xu. This work was supported by the National Natural Science Foundation of China (62001044, 62201013), the Basic Research Program of China (JCKY2021210B059), the Equipment Advance Research Field Foundation (315067206), and the Fund of State Key Laboratory of Information Photonics and Optical Communications (IPOC2021ZT02).